# A Full-Density Approach to Simulating Random Iteration Equations with Applications

### 29 Mar 2026 v2

@@ Abstract
The goal of this study is to introduce a unified computational framework for simulating random iteration equations (RIE), understood as iteration equations containing random variables. The novelty of this work is that full probability densities of the state vectors are propagated stepwise through the iterations avoiding the need of repetitive pathwise Monte Carlo simulations of the iteration equation. The presentation of the methodology is conceptually efficient based on recent work on static random equations and intentionally accessible. The technical requirements on the RIE are minimal based on the previous work, allowing for potential nonlinearities, discontinuities and stochasticities in the transfer function, as well as nonstandard densities and diffusion processes. As results, illustrative applications of random and stochastic differential equation simulations, a novel full-density gradient descent method (FDGD) for global optimization under uncertainty and examples of chaotic mappings are presented in order to demonstrate the breadth of the utility of this framework. In total, the character of the presentation is explorative and encourages new applications and theoretical studies.
@@ 2.1 General Formulation
leading to $\boldsymbol{B}$ a random variable vector centered around zero and with small standard deviations compared to other standard deviations and the required solution vector $\boldsymbol{x}^{(n+1)}$ for each iteration. The rather technical minimal assumptions about $\boldsymbol{T}$ are inherited directly from the integrability assumptions about $\boldsymbol{M}$ presented in [6] and are typically only violated in constructed limit cases. In total, this leads to
@@
$\mathcal{L}_{\boldsymbol{x}^{(n+1)}}(\boldsymbol{0}|\boldsymbol{x})=\int\limits_{\mathbb{R}^{K-R}}\int\limits_{\mathbb{R}^{R}}f_{\boldsymbol{B}}(\boldsymbol{x}-\boldsymbol{T}(\boldsymbol{s}_{1},\boldsymbol{s}_{2}))\cdot f_{\boldsymbol{x}^{(n)}}(\boldsymbol{s}_{1})\cdot f_{\boldsymbol{C}}(\boldsymbol{s}_{2})\;\text{d}\boldsymbol{s}_{1}\,\text{d}\boldsymbol{s}_{2}\;.$ (5)
@@ 2.3 Discretization of Dynamical Systems with Stochasticity as Random Iteration Equation
A major observation is that when utilizing this discretization scheme, we can also simulate the evolution of probability density functions of SDEs with random coefficients utilizing the Euler-Maruyama algorithm [4, 5, 10] by introducing new terms
@@
with the diffusion coefficient function $\boldsymbol{B}(t_{n},\boldsymbol{x}^{(n)},\boldsymbol{D}^{(n)})$, which may depend on a random variable vector $\boldsymbol{D}^{(n)}$, and a Wiener process $\boldsymbol{W}_{t}$ with a discrete increment $\Delta\boldsymbol{W}^{(n)}=\boldsymbol{W}_{t_{n+1}}-\boldsymbol{W}_{t_{n}}$ as normally distributed random variables with zero mean and variances $\Delta t$. This leads to the general transfer function for SDEs
@@ 4 Discussion
There are several dimensions to further develop the application of the framework: a) more efficient numerical schemes for sampling the density functions in the Monte Carlo integration than, e.g. acceptance rejection sampling, b) more efficient use of the definition of the state space grid in the simulation, including possibly adaptive grids, c) better discretization approaches in the derivation of the random iteration equations (e.g. for continuous dynamical systems from Euler algorithm to Runge Kutta algorithms, e.g. [22]) and a detailed investigation of their convergence rates, d) extension of the optimization approaches, e.g. higher order iterative schemes such as full-density Newton optimization, etc., e) more challenging parameter density functions apart from normal distributions and a broader (also discontinuous) selection of the transfer functions, f) going from $R=2$ to higher dimensional state spaces and/or random variable vectors, e.g. investigating the curse of dimensionality for this approach, and finally, g) extension to other interesting application areas where iteration equations can play a productive role in understanding dynamical phenomena and contain random variable densities as a new degree of freedom in the modeling process.

### 02 Apr 2026 v3

@@ Abstract
The goal of this study is to introduce a unified computational framework for simulating random iteration equations (RIE), understood as iteration equations containing random variables. The novelty of this work is that full probability densities of the state vectors are propagated stepwise through the iterations avoiding the need of repetitive pathwise Monte Carlo simulations of the iteration equation. The presentation of the methodology is conceptually efficient based on recent work on static random equations and intentionally accessible. The technical requirements on the RIE are mild based on the previous work, allowing for potential nonlinearities, discontinuities and stochasticities in the transfer function, as well as nonstandard densities and diffusion processes. As results, illustrative applications of random and stochastic differential equation simulations, a novel full-density gradient descent method (FDGD) for global optimization under uncertainty and examples of chaotic mappings are presented in order to demonstrate the breadth of the utility of this framework. In total, the character of the presentation is explorative and encourages new applications and theoretical studies.
@@ 2.1 General Formulation
leading to $\boldsymbol{B}$ a random variable vector centered around zero and with small standard deviations compared to other standard deviations and the solution vector $\boldsymbol{x}^{(n+1)}$ for each iteration. The mild technical assumptions about $\boldsymbol{T}$ are inherited directly from the integrability assumptions about $\boldsymbol{M}$ presented in [6] and are typically only violated in constructed limit cases. In total, this leads to
@@
$\mathcal{L}_{\boldsymbol{x}^{(n+1)}}(\boldsymbol{0}|\boldsymbol{x})=\int\limits_{\mathbb{R}^{K}}\int\limits_{\mathbb{R}^{R}}f_{\boldsymbol{B}}(\boldsymbol{x}-\boldsymbol{T}(\boldsymbol{s}_{1},\boldsymbol{s}_{2}))\cdot f_{\boldsymbol{x}^{(n)}}(\boldsymbol{s}_{1})\cdot f_{\boldsymbol{C}}(\boldsymbol{s}_{2})\;\text{d}\boldsymbol{s}_{1}\,\text{d}\boldsymbol{s}_{2}\;.$ (5)
@@
Rewriting Equation 5 into
$\pi_{\boldsymbol{x}^{(n+1)}}(\boldsymbol{x})\propto\int\limits_{\mathbb{R}^{R}}\underbrace{\left(\;\int\limits_{\mathbb{R}^{K}}f_{\boldsymbol{B}}(\boldsymbol{x}-\boldsymbol{T}(\boldsymbol{s}_{1},\boldsymbol{s}_{2}))\cdot f_{\boldsymbol{C}}(\boldsymbol{s}_{2})\,\text{d}\boldsymbol{s}_{2}\right)}_{=:k(\boldsymbol{x},\boldsymbol{s}_{1})}\cdot\;\pi_{\boldsymbol{x}^{(n)}}(\boldsymbol{s}_{1})\;\text{d}\boldsymbol{s}_{1}\;,$ (7)
the RIE iteration can be reframed as an integral linear operator acting on the posterior densities with kernel $k(\boldsymbol{x},\boldsymbol{s}_{1})$, which we denote as the linear RIE operator. We emphasize that this reformulation is in general possible including cases with nonlinearities, discontinuities or untypical parameter probability densities in the transfer function $\boldsymbol{T}$. Spectral investigation for specific kernels may lead to further insights into the dynamic properties of the density propagation for selected RIEs.
@@ 2.2 Numerical Implementation
The major computational compromise must be found between the number of samples $P$, the standard deviations of $\boldsymbol{B}$ and the accepted noise level in the final density by utilizing Monte Carlo integration. For example, the more concentrated $\boldsymbol{B}$ is around zero (what is desirable since too large standard deviations introduce a blur on the posterior density or even slight diffusion over a large number of iterations), the more samples inserted in $f_{B_{r}}$ are typically not contributing to the final posterior density increasing the noise level.
@@
It is pointed out that utilizing the perspective of the linear RIE operator combined with a linear decomposition of the posterior density function as linear combination of basis functions, much more efficient numerical schemes are possible than the presented Monte Carlo algorithm. In this case the coefficients of the linear combination are updated in each iteration by multiplying them with a propagation matrix. This perspective can also be utilized as an efficient method to incorporate spatial discretization directly into the algorithm by (effectively) pixel basis functions.
@@ 2.3 Discretization of Dynamical Systems with Stochasticity as Random Iteration Equation
A major observation is that when utilizing this discretization scheme, we can also simulate the evolution of probability density functions of SDEs with random coefficients
$\text{d}\boldsymbol{x}=\boldsymbol{F}(t,\boldsymbol{x},\boldsymbol{C})\,\text{d}t+\boldsymbol{B}(t,\boldsymbol{x},\boldsymbol{D})\,\text{d}\boldsymbol{W}_{t}\;,$ (12)
with the drift coefficent function $\boldsymbol{F}$, the diffusion coefficient function $\boldsymbol{B}$ and a Wiener process $\boldsymbol{W}_{t}$, utilizing the Euler-Maruyama algorithm [4, 5, 10] by introducing new terms
@@
with the coefficent random variable vectors $\boldsymbol{C}^{(n)}$ and $\boldsymbol{D}^{(n)}$, and the discrete Wiener process increment $\Delta\boldsymbol{W}^{(n)}=\boldsymbol{W}_{t_{n+1}}-\boldsymbol{W}_{t_{n}}$ as normally distributed random variables with zero mean and variances $\Delta t$. This leads to the general transfer function for SDEs by renaming the random variables
@@ 4 Discussion
There are several dimensions to further develop the application of the framework: a) more efficient numerical schemes for sampling the density functions in the Monte Carlo integration than, e.g. acceptance rejection sampling, as well as utilizing the linear RIE operator combined with a linear decomposition of the posterior density functions, b) more efficient use of the definition of the state space grid in the simulation, including possibly adaptive grids, c) better temporal discretization approaches in the derivation of the random iteration equations (e.g. for continuous dynamical systems from Euler algorithm to Runge Kutta algorithms, e.g. [22]) and a detailed investigation of their convergence rates, d) extension of the optimization approaches, e.g. higher order iterative schemes such as full-density Newton optimization, etc., e) more challenging parameter density functions apart from normal distributions and a broader (also discontinuous) selection of the transfer functions, f) going from $R=2$ to higher dimensional state spaces and/or random variable vectors, e.g. investigating the curse of dimensionality for this approach, and finally, g) extension to other interesting application areas where iteration equations can play a productive role in understanding dynamical phenomena and contain random variable densities as a new degree of freedom in the modeling process.